OpenStudy (anonymous):

write an equation in standard form of the line passes through (0,3) (2,-3)

OpenStudy (anonymous):

1) Find the slope m = -6/2 = -3 y = -3 + 3 Note the last number is the y intercept, which was ginev in this problem

OpenStudy (anonymous):

for the answer to the first question on the very top the standard form is y=-3x+3

OpenStudy (anonymous):

Yes, @samsterz is right I forgot the x, sorry

OpenStudy (anonymous):

ok thank you!!